OpenStudy (anonymous):

Original Post: Y = squarerootx Find the derivate of the function using the definition of derivate Y=squarerootx

OpenStudy (tkhunny):

Wow!

OpenStudy (mathmale):

It appears to me that this question is both legitimate and complete. Using a particular method, "the definition of the derivative," @lizzy92 is to find the derivative of \[y=\sqrt{x}\]

OpenStudy (mathmale):

The definition of the derivative of f(x) is as follows:\[f '(x) =limit~(as~h~approaches~zero)~of ~\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (mathmale):

Here, \[f(x)=\sqrt{x}~and~f(x+h)=\sqrt{x+h}\]

OpenStudy (mathmale):

lizzy: what further help do you need at this point?

OpenStudy (mathmale):

@lizzy92: ?

OpenStudy (aum):

\[ \lim_{h \rightarrow 0}\frac{\sqrt{x+h}-\sqrt{x}}{h} = ? \]Multiply top and bottom by the conjugate of the numerator \(\sqrt{x+h}+\sqrt{x}\) and simplify and then take the limit.

OpenStudy (aum):

Make use of the identity: \((a-b)(a+b) = a^2 - b^2\) after you multiply the numerator by its conjugate.

OpenStudy (anonymous):

Ok thank you guys! :)

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Thank you @mathmale :)